Car batteries are often rated in amp-hours. This rating indicates the quantity of charge that can pass through the battery before the batterys chemical energy must be restored by the cars generator or some other source. The amp-hour unit is convenient for simple calculations. For example, a 60-amp-hour battery can supply a current of 1.0 A for 60 h (or 60 A for 1.0 h) before going dead. Find the energy stored in the following. (a) a 54 A-h, 14 V car battery (b) a 1.4 A-h, 1.4 V flashlight battery
Answer #1

Solution :-

a) Energy stored = V*i*time = 54×14×3600 = 2721600J

b)
Energy stored =v*i*time=1.4*1.4*3600 = 7056J